Famous Productions performs London shows. The average show sells 1,300 tickets at $60 per ticket. There are 145 shows per yeaRequirement 1. Compute revenue and variable costs for each show. Select the formula and enter the amounts to compute sales reSelect the formula and enter the amounts to compute variable costs for each show. Compute the variable costs per show for eacRequirement 2. Use the equation approach to compute the number of shows Famous Productions must perform each year to break evNow use the information given and the required sales in dollars computed in the previous step to determine the required numbeRequirement 4. Prepare Famous Productionss contribution margin income statement for 145 shows performed in 2018. Report only
Answer #1

1.Sales Revenue per show = Number of Tickets*Price per ticket

= 1300*60

= $78000

Variable cost per show

Cost per unit*Number of units

Cost of program = 1300*8 = $10400

Cost of performers = 65*340 = $22100

Total variable cost = $32500

2.Sales Revenue – Variable costs – Fixed costs = Target Profit

Number of shows to break even = 728000/45500

= 16 shows

3.Number of shows required = (5687500+728000)/45500

= 144.96 shows

i.e. 145 shows

Yes, realistic

Contribution Margin Income Statement

Sales Revenue

        11,310,000

Less: Variable costs

          4,712,500

Contribution Margin

          6,597,500

Less: Fixed costs

             728,000

Net Operating Income

          5,869,500